Given f(x) =[1/3 + x/66] , then `sigma x=1to66 f(x) ` is(ans is 23)

Grade:12th pass

1 Answers

Sanyu daver
6 years ago
Here f(x) can be written as f(x)= 1/3 +[x/66] So sigma f(x) from 1 to 66 will be 1/3 ×66 (as 1/3 will be added 66 times)+ greatest integer upto x= 65 will be 0 and x=66 will be 1 so 1/3×66+1 = 22+1=23